# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ III Locally Differentially Private Distributed Aggregative Optimization Algorithm
IV Optimization Accuracy and Convergence Rate Analysis
@@
For the sake of notational simplicity, we denote the lowest decaying rates of DP-noise variances among all agents as $\varsigma_{x}=\min_{i\in[m]}\{\varsigma_{i,x}\}$, $\varsigma_{y}=\min_{i\in[m]}\{\varsigma_{i,y}\}$, and $\varsigma_{z}=\min_{i\in[m]}\{\varsigma_{i,z}\}$, and the optimization variable of all agents as $\boldsymbol{x}^{t}=\mathrm{col}(x_{1}^{t},\cdots,x_{m}^{t})$.
@@
(i) if $F(\boldsymbol{x})$ is strongly convex with the rates of DP-noise variances satisfying $\varsigma_{x}>\max\{v_{z}-\varsigma_{z},v_{y}-\varsigma_{y},\frac{v_{x}}{2}\}$, $\varsigma_{y}>v_{y}-\frac{1}{2}$, and $\varsigma_{z}>v_{z}-\frac{1}{2}$, then we have
$\mathbb{E}[\|\boldsymbol{x}^{T}-\boldsymbol{x}^{*}\|^{2}]\leq\mathcal{O}(T^{-\beta}),\vspace{-0.2em}$ (4)
with $\beta=\min\{2\varsigma_{x}-v_{x},\frac{1}{2}-v_{y}+\varsigma_{y},\frac{1}{2}-v_{z}+\varsigma_{z}\}$.
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]\leq\mathcal{O}(T^{-(1-v_{x})}).\vspace{-0.2em}$ (5)
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[\|\nabla F(\boldsymbol{x}^{t})\|^{2}]\leq\mathcal{O}(T^{-(1-v_{x})}).$ (6)
@@
Corollary 1 proves that the convergence rate of our algorithm is arbitrarily close to $\mathcal{O}(T^{-0.5})$ for nonconvex/convex/strongly convex objective functions even in the presence of Laplace noise. This is consistent with the existing result in [22], which established a convergence rate of $\mathcal{O}(T^{-0.5})$ for the special case of convex $F(\boldsymbol{x})$ and noise-free aggregative function $g(\boldsymbol{x})$.
@@ V Privacy analysis
(Sensitivity [48]) We denote the adjacent relationship between $\mathcal{D}_{i}^{t}$ and $\mathcal{D}_{i}^{\prime t}$ by $\text{Adj}(\mathcal{D}_{i}^{t},\mathcal{D}_{i}^{\prime t})$. The sensitivity of each agent $i$’s implementation $\mathcal{A}_{i}$ at time $t$ is defined as
$\Delta_{i}^{t+1}=\max_{\text{Adj}(\mathcal{D}_{i}^{t},\mathcal{D}_{i}^{\prime t})}\|\mathcal{A}_{i}(\mathcal{D}_{i}^{t},\theta_{-i}^{t})-\mathcal{A}_{i}(\mathcal{D}_{i}^{\prime t},\theta_{-i}^{t})\|_{1},\vspace{-0.3em}$
@@
According to Definition 3, under Algorithm 1, agent $i$’s implementation involves three sensitivities: $\Delta_{i,x}^{t}$, $\Delta_{i,y}^{t}$, and $\Delta_{i,z}^{t}$, which correspond to the three shared variables $\boldsymbol{x}_{i}^{t}+\boldsymbol{\vartheta}_{i}^{t}$, $y_{i}^{t}+\chi_{i}^{t}$, and $z_{i}^{t}+\zeta_{i}^{t}$, respectively. With this understanding, we present the following lemma, which delineates a sufficient condition for $\epsilon_{i}$-LDP over any time horizon $T\in\mathbb{N}^{+}$:
@@
For any given $T\in\mathbb{N}^{+}$ (which includes the case of $T\rightarrow\infty$), if agent $i$ injects to each of its shared variables $\boldsymbol{x}_{i}^{t}$, $y_{i}^{t}$, and $z_{i}^{t}$ at each iteration $t\in\{1,\cdots,T\}$ noise vectors $\boldsymbol{\vartheta}_{i}^{t}$, $\chi_{i}^{t}$, and $\zeta_{i}^{t}$ consisting of $n$, $r$, and $r$ independent Laplace noises with parameters $\nu_{i,x}^{t}$, $\nu_{i,y}^{t}$, and $\nu_{i,z}^{t}$, respectively, then agent $i$’s implementation $\mathcal{A}_{i}$ is $\epsilon_{i}$-locally differentially private with the cumulative privacy budget from $t=1$ to $t=T$ upper bounded by $\sum_{t=1}^{T}(\frac{\Delta_{i,x}^{t}}{\nu_{i,x}^{t}}+\frac{\Delta_{i,y}^{t}}{\nu_{i,y}^{t}}+\frac{\Delta_{i,z}^{t}}{\nu_{i,z}^{t}})$.
@@
Without loss of generality, we consider adjacent datasets $\mathcal{D}_{i,f}^{t}$ (resp., $\mathcal{D}_{i,g}^{t}$) and $\mathcal{D}_{i,f}^{\prime t}$ (resp., $\mathcal{D}_{i,g}^{\prime t}$) that differ in the $k$th element. For the sake of clarity, the variables corresponding to $\mathcal{D}_{i,f}^{t}$ and $\mathcal{D}_{i,g}^{t}$ are denoted as $\boldsymbol{x}_{i}^{t}$, $y_{i}^{t}$, and $z_{i}^{t}$. Similarly, the variables corresponding to $\mathcal{D}_{i,f}^{\prime t}$ and $\mathcal{D}_{i,g}^{\prime t}$ are denoted as $\boldsymbol{x}_{i}^{\prime t}$, $y_{i}^{\prime t}$, and $z_{i}^{\prime t}$.
@@
Under Assumptions 2-5 and the conditions in Theorem 1, $\boldsymbol{x}^{t}$ (resp., $F(\boldsymbol{x}^{t})$ in the general convex case and $\nabla F(\boldsymbol{x}^{t})$ in the nonconvex case) in Algorithm 1 converges in mean-square to the optimal solution $\boldsymbol{x}^{*}$ to problem (1) (resp., in mean to $F(\boldsymbol{x}^{*})$ and in mean-square to zero). Furthermore,
(i) for any $T\in\mathbb{N}^{+}$ (which includes the case of $T\rightarrow\infty$), agent $i$’s implementation of Algorithm 1 is locally differentially private with a cumulative privacy budget bounded by $\epsilon_{i}=\epsilon_{i,x}+\epsilon_{i,y}+\epsilon_{i,z}$, where $\epsilon_{i,x}$, $\epsilon_{i,y}$, and $\epsilon_{i,z}$ are given by $\epsilon_{i,x}\leq\sum_{t=1}^{T}\frac{\sqrt{2}C_{x}}{\sigma_{i,x}(t+1)^{1+v_{x}-v_{z}-\varsigma_{i,x}}}$, $\epsilon_{i,y}\leq\sum_{t=1}^{T}\frac{\sqrt{2}C_{y}}{\sigma_{i,y}(t+1)^{1+v_{y}-\varsigma_{i,y}}}$, and $\epsilon_{i,z}\leq\sum_{t=1}^{T}\frac{\sqrt{2}C_{z}}{\sigma_{i,z}(t+1)^{1+v_{z}-\varsigma_{i,z}}}$ with the positive constants $C_{x}$, $C_{y}$, and $C_{z}$ given in (18), (16), and (19), respectively.
@@
According to Definition 3, we have $\boldsymbol{x}_{j}^{t}+\boldsymbol{\vartheta}_{j}^{t}=\boldsymbol{x}_{j}^{\prime t}+\boldsymbol{\vartheta}_{j}^{\prime t}$, $y_{j}^{t}+\chi_{j}^{t}=y_{j}^{\prime t}+\chi_{j}^{\prime t}$, and $z_{j}^{t}+\zeta_{j}^{t}=z_{j}^{\prime t}+\zeta_{j}^{\prime t}$ for all $t\in\mathbb{N}$ and $j\in{\mathcal{N}_{i}}$. Since we assume that only the $k$th data point is different between $\mathcal{D}_{i,f}^{t}$ (resp. $\mathcal{D}_{i,g}^{t}$) and $\mathcal{D}_{i,f}^{\prime t}$ (resp. $\mathcal{D}_{i,g}^{\prime t}$), we have $\boldsymbol{x}_{i}^{p}=\boldsymbol{x}_{i}^{\prime p}$, $y_{i}^{p}=y_{i}^{\prime p}$, and $z_{i}^{p}=z_{i}^{\prime p}$ for all $p\leq k\leq t$. However, since the difference in loss functions kicks in at iteration $k$, i.e., $h(x,y;\varphi_{i}^{k})\neq h(x,y;\varphi_{i}^{\prime k})$ and $l(x;\xi_{i}^{k})\neq l(x;\xi_{i}^{\prime k})$, we have $\boldsymbol{x}_{i}^{t}\neq\boldsymbol{x}_{i}^{\prime t}$, $y_{i}^{t}\neq y_{i}^{\prime t}$, and $z_{i}^{t}\neq z_{i}^{\prime t}$ for all $t>k$. Hence, for the dynamics of $y_{i}^{t}$ in Algorithm 1, we have
@@
$\textstyle\quad+\lambda_{y}^{t}(g_{i}^{t}(R_{i}\boldsymbol{x}_{i}^{t})-g_{i}^{\prime t}(R_{i}\boldsymbol{x}_{i}^{\prime t}))\|_{1},$
@@
By defining $\bar{w}\triangleq\min_{i\in[m]}\{|w_{ii}|\}$ and using the relationship $\xi_{i}^{p}=\xi_{i}^{\prime p}$ for any $p\neq k$, the sensitivity $\Delta_{i,y}^{t+1}$ satisfies
$\textstyle\Delta_{i,y}^{t+1}$ $\textstyle\leq(1-\bar{w})\Delta_{i,y}^{t}+\frac{\lambda_{y}^{t}}{t+1}\|l(R_{i}\boldsymbol{x}_{i}^{t};\xi_{i}^{k})-l(R_{i}\boldsymbol{x}_{i}^{\prime t};\xi_{i}^{\prime k})\|_{1}$
$\textstyle\quad+\frac{\sqrt{r}\lambda_{y}^{t}}{t+1}\sum_{p\neq k}^{t}\|l(R_{i}\boldsymbol{x}_{i}^{t};\xi_{i}^{p})-l(R_{i}\boldsymbol{x}_{i}^{\prime t};\xi_{i}^{p})\|_{2},$
@@
$\Delta_{i,y}^{t+1}\leq(1-\bar{w})\Delta_{i,y}^{t}+L_{l}\sqrt{r}\lambda_{y}^{t}\Delta_{i,x}^{t}+\frac{2d_{l}\lambda_{y}^{t}}{t+1},$ (7)
where in the derivation we have used the fact that the optimization variable $\boldsymbol{x}_{i}^{t}$ is always constrained in the compact set $\Omega$, i.e., $\max_{i\in[m]}\sup_{\boldsymbol{x}_{i}\in\Omega}\{\|l(R_{i}\boldsymbol{x}_{i};\xi_{i})\|_{1}\}\leq d_{l}$ holds for some $d_{l}>0$.
Furthermore, we use the projection inequality from Lemma 1 in [63] and the dynamics of $\boldsymbol{x}_{i}^{t}$ to obtain
$\textstyle\|\boldsymbol{x}_{i}^{t+1}-\boldsymbol{x}_{i}^{\prime t+1}\|_{1}\leq(1-\bar{w})\|\boldsymbol{x}_{i}^{t}-\boldsymbol{x}_{i}^{\prime t}\|_{1}$
$\textstyle+\frac{\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla_{x}h(R_{i}\boldsymbol{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{p})-\nabla_{x}h(R_{i}\boldsymbol{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{\prime p})\|_{1}$
$\textstyle+\frac{\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla l(R_{i}\boldsymbol{x}_{i}^{t};\xi_{i}^{p})\tilde{z}_{i}^{t}-\nabla l(R_{i}\boldsymbol{x}_{i}^{\prime t};\xi_{i}^{\prime p})\tilde{z}_{i}^{\prime t}\|_{1}.$ (8)
@@
$\textstyle\frac{\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla_{x}h(R_{i}\boldsymbol{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{p})-\nabla_{x}h(R_{i}\boldsymbol{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{\prime p})\|_{1}$
$\textstyle\leq\frac{\sqrt{n_{i}}\lambda_{x}^{t}}{t+1}\sum_{p\neq k}^{t}\|\nabla_{x}h(R_{i}\boldsymbol{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{p})-\nabla_{x}h(R_{i}\boldsymbol{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{p})\|_{2}$
$\textstyle\quad+\frac{\sqrt{n_{i}}\lambda_{x}^{t}}{t+1}\|\nabla_{x}h(R_{i}\boldsymbol{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{k})-\nabla_{x}h(R_{i}\boldsymbol{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{\prime k})\|_{2}$
$\textstyle\leq\frac{\sqrt{n_{i}}\bar{L}_{h}\lambda_{x}^{t}t}{t+1}(\|\boldsymbol{x}_{i}^{t}-\boldsymbol{x}_{i}^{\prime t}\|_{2}\!+\!\|\tilde{y}_{i}^{t}-\tilde{y}_{i}^{\prime t}\|_{2})\!+\!\frac{2\sqrt{n_{i}}L_{h}\lambda_{x}^{t}}{t+1}$
@@
where we have used the relationship $\|\tilde{y}_{i}^{t}-\tilde{y}_{i}^{\prime t}\|_{2}\leq\frac{1}{\lambda_{y}^{t}}(\|y_{i}^{t+1}-y_{i}^{\prime t+1}\|_{1}+\|y_{i}^{t}-y_{i}^{\prime t}\|_{1})$ in the last inequality.

### v3

@@ III Locally Differentially Private Distributed Aggregative Optimization Algorithm
Remark 2.
In Algorithm 1, although agent $i$ estimates the optimization variables of other agents, it cannot reconstruct the private data of others. Specifically, for any finite number of iterations, i.e., $T\!<\!\infty$, each agent $i$ does not have access to the exact optimization variables of others and only receives noise-corrupted variables from neighboring agents, which prevent accurate recovery of the exact optimization variables or inference of the private data of others. Furthermore, in the asymptotic case, i.e., $T\!=\!\infty$, each agent $i$ can indeed accurately know the optimal solutions of other agents. However, the optimal solution, i.e., $\mathbf{x}_{i}^{\infty}\!=\!\boldsymbol{x}^{*}$, is only a single point in the decision space $\mathbb{R}^{n}$ and is insufficient to reconstruct the objective functions or the underlying private data of other agents. In fact, revealing such a single point $\boldsymbol{x}^{*}$ corresponds to disclosing an event of measure zero with respect to the data-generating distribution, which does not affect the probabilities of outputs and therefore does not violate the $\epsilon_{i}$-LDP guarantees established based on output probabilities.
Remark 3.
Our algorithm can, in principle, be extended to directed graphs by leveraging the push-sum method [28, 49, 42], which introduces normalization factors to compensate for network imbalances. Since the updates of these normalization factors are independent of the optimization variables, integrating the push-sum mechanism with Algorithm 1 is conceptually feasible. However, rigorously incorporating these normalization factors would require introducing additional communication variables and substantially more involved technical derivations, leading to a significantly lengthier convergence analysis and privacy analysis. We leave a systematic treatment of this extension to future work.
Remark 4.
Our work has many natural applications in practical control systems. For example, [65] shows that in DC microgrid control systems, the objective of voltage regulation under current-sharing and economic operation requirements can be naturally formulated as a distributed aggregative optimization problem, thereby explicitly casting the control design task within an aggregative optimization framework (see Eq. (3) in [65]). Similarly, [23] shows that multi-agent formation control problems, e.g., the target surrounding problem, can be formulated as a distributed aggregative optimization problem, making Algorithm 1 applicable to these types of control tasks.
IV Convergence Accuracy and Rate Analysis
@@
For the sake of notational simplicity, we denote the lowest decaying rates of DP-noise variances among agents as $\varsigma_{x}\!=\!\min_{i\in[m]}\{\varsigma_{i,x}\}$, $\varsigma_{y}\!=\!\min_{i\in[m]}\{\varsigma_{i,y}\}$, and $\varsigma_{z}\!=\!\min_{i\in[m]}\{\varsigma_{i,z}\}$.
@@
(i) if $F(\boldsymbol{x})$ is strongly convex with the rates of DP-noise variances satisfying $\varsigma_{x}>\max\{v_{y}-\varsigma_{y},v_{z}-\varsigma_{z},\frac{v_{x}}{2}\}$, $\varsigma_{y}>v_{y}-\frac{1}{2}$, and $\varsigma_{z}>v_{z}-\frac{1}{2}$, then we have
$\mathbb{E}[\|\mathbf{x}_{i}^{T}-\boldsymbol{x}^{*}\|^{2}]\leq\mathcal{O}(T^{-\beta}),$ (5)
with $\beta=\min\{2\varsigma_{x}-v_{x},\frac{1}{2}-v_{y}+\varsigma_{y},\frac{1}{2}-v_{z}+\varsigma_{z}\};$
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[F(\mathbf{x}_{i}^{t})-F(\boldsymbol{x}^{*})]\leq\mathcal{O}(T^{-(1-v_{x})});$ (6)
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[\|\nabla F(\mathbf{x}_{i}^{t})\|^{2}]\leq\mathcal{O}(T^{-(1-v_{x})}).$ (7)
@@
Corollary 1 proves that the convergence rate of our algorithm is arbitrarily close to $\mathcal{O}(T^{-0.5})$ for nonconvex/convex/strongly convex objective functions even in the presence of Laplace noise. This is consistent with the existing result in [11], which established a convergence rate of $\mathcal{O}(T^{-0.5})$ for the special case of convex $F(\boldsymbol{x})$ and noise-free aggregative function $g(\boldsymbol{x})$. Moreover, Corollary 1 provides a simple way to select parameters that satisfy the conditions in Theorem 1. Specifically, it implies that for any $\delta\in(0,\frac{1}{14})$, the parameters $v_{x}$, $v_{y}$, $v_{z}$, $\varsigma_{x}$, $\varsigma_{y}$, and $\varsigma_{z}$, chosen according to the specified rules, will satisfy the conditions in Theorem 1. Hence, in real-world applications, we do not need to tune six parameters $v_{x}$, $v_{y}$, $v_{z}$, $\varsigma_{x}$, $\varsigma_{y}$, and $\varsigma_{z}$, but instead only need to choose one parameter $\delta$.
Corollary 2.
Under Assumption 1, for any $T\in\mathbb{N}^{+}$ and $i\in[m]$, the communication and computational complexities of Algorithm 1 are on the order of $\mathcal{O}((2r+n)d_{i}T)$ and $\mathcal{O}(n_{i}rT^{2})$, respectively, where $d_{i}$ denotes the number of agent $i$’s neighbors.
Proof.
See Appendix C. ∎
Compared with the conventional distributed aggregative optimization algorithms that do not consider privacy in, e.g., [35, 36], where the dimension of exchanged messages is $2r$, Algorithm 1 requires exchanging messages of dimension $2r+n$. This increased communication overhead arises from the imposed LDP constraints. Specifically, the noise added to satisfy the LDP requirements inevitably reduces the amount of useful information conveyed in each shared variable. As a result, more variable exchanges are required to transmit sufficient information to ensure accurate convergence, compared with the conventional distributed aggregative optimization algorithms that ignore privacy. We regard this increased communication overhead as a necessary price for achieving accurate convergence under rigorous $\epsilon_{i}$-LDP.
In addition, we note that using historical data to compute gradients increases the computational complexity compared with the strategy of using one data-point per iteration. However, for continuous loss functions, our prior work [14, Section III-B] shows that, by employing Lagrange interpolation, the increased computational complexity at each iteration can be made independent of the iteration number. As a result, the computational complexity of Algorithm 1 over $T$ iterations can be reduced from $\mathcal{O}(n_{i}rT^{2})$ to $\mathcal{O}(n_{i}rT)$, which is consistent with the order of computational complexity of noise-free distributed aggregative optimization algorithms in, e.g., [36].
@@ V Privacy analysis
(Sensitivity [51]) We denote the adjacency relationship between $\mathcal{D}_{i}^{t}$ and $\mathcal{D}_{i}^{\prime t}$ by $\text{Adj}(\mathcal{D}_{i}^{t},\mathcal{D}_{i}^{\prime t})$. The sensitivity of each agent $i$’s implementation $\mathcal{A}_{i}$ at time $t$ is defined as
$\Delta_{i}^{t+1}=\max_{\text{Adj}(\mathcal{D}_{i}^{t},\mathcal{D}_{i}^{\prime t})}\|\mathcal{A}_{i}(\mathcal{D}_{i}^{t},\theta_{-i}^{t})-\mathcal{A}_{i}(\mathcal{D}_{i}^{\prime t},\theta_{-i}^{t})\|_{1},\vskip-3.00003pt$
@@
According to Definition 3, under Algorithm 1, agent $i$’s implementation involves three sensitivities: $\Delta_{i,x}^{t}$, $\Delta_{i,y}^{t}$, and $\Delta_{i,z}^{t}$, which correspond to the three shared variables $\mathbf{x}_{i}^{t}+\boldsymbol{\vartheta}_{i}^{t}$, $y_{i}^{t}+\chi_{i}^{t}$, and $z_{i}^{t}+\zeta_{i}^{t}$, respectively. With this understanding, we present the following lemma, which delineates a sufficient condition for $\epsilon_{i}$-LDP over any time horizon $T\in\mathbb{N}^{+}$:
@@
For any given $T\in\mathbb{N}^{+}$ (which includes the case of $T\rightarrow\infty$), if agent $i$ injects to each of its shared variables $\mathbf{x}_{i}^{t}$, $y_{i}^{t}$, and $z_{i}^{t}$ at each iteration $t\in\{1,\cdots,T\}$ noise vectors $\boldsymbol{\vartheta}_{i}^{t}$, $\chi_{i}^{t}$, and $\zeta_{i}^{t}$ consisting of $n$, $r$, and $r$ independent Laplace noises with parameters $\nu_{i,x}^{t}$, $\nu_{i,y}^{t}$, and $\nu_{i,z}^{t}$, respectively, then agent $i$’s implementation $\mathcal{A}_{i}$ is $\epsilon_{i}$-locally differentially private with the cumulative privacy budget from $t=1$ to $t=T$ upper bounded by $\sum_{t=1}^{T}(\frac{\Delta_{i,x}^{t}}{\nu_{i,x}^{t}}+\frac{\Delta_{i,y}^{t}}{\nu_{i,y}^{t}}+\frac{\Delta_{i,z}^{t}}{\nu_{i,z}^{t}})$.
@@
Without loss of generality, we consider adjacent datasets $\mathcal{D}_{i,f}^{t}$ (resp. $\mathcal{D}_{i,g}^{t}$) and $\mathcal{D}_{i,f}^{\prime t}$ (resp. $\mathcal{D}_{i,g}^{\prime t}$) that differ in the $k$th element. For the sake of clarity, the variables corresponding to $\mathcal{D}_{i,f}^{t}$ and $\mathcal{D}_{i,g}^{t}$ are denoted as $\mathbf{x}_{i}^{t}$, $y_{i}^{t}$, and $z_{i}^{t}$. Similarly, the variables corresponding to $\mathcal{D}_{i,f}^{\prime t}$ and $\mathcal{D}_{i,g}^{\prime t}$ are denoted as $\mathbf{x}_{i}^{\prime t}$, $y_{i}^{\prime t}$, and $z_{i}^{\prime t}$.
@@
Under Assumptions 2-5 and the conditions in Theorem 1, $\mathbf{x}_{i}^{t}$ (resp. $F(\mathbf{x}_{i}^{t})$ in the general convex case and $\nabla F(\mathbf{x}_{i}^{t})$ in the nonconvex case) in Algorithm 1 converges in mean-square to the optimal solution $\boldsymbol{x}^{*}$ to problem (1) (resp. in mean to $F(\boldsymbol{x}^{*})$ and in mean-square to zero). Furthermore,
(i) for any $T\in\mathbb{N}^{+}$ (which includes the case of $T\rightarrow\infty$), agent $i$’s implementation of Algorithm 1 is locally differentially private with a cumulative privacy budget bounded by $\epsilon_{i}=\epsilon_{i,x}+\epsilon_{i,y}+\epsilon_{i,z}$, where $\epsilon_{i,x}$, $\epsilon_{i,y}$, and $\epsilon_{i,z}$ are given by $\epsilon_{i,x}\leq\sum_{t=1}^{T}\frac{\sqrt{2}C_{i,x}}{\sigma_{i,x}(t+1)^{1+v_{x}-v_{z}-\varsigma_{i,x}}}$, $\epsilon_{i,y}\leq\sum_{t=1}^{T}\frac{\sqrt{2}C_{i,y}}{\sigma_{i,y}(t+1)^{1+v_{y}-\varsigma_{i,y}}}$, and $\epsilon_{i,z}\leq\sum_{t=1}^{T}\frac{\sqrt{2}C_{i,z}}{\sigma_{i,z}(t+1)^{1+v_{z}-\varsigma_{i,z}}}$ with the positive constants $C_{i,x}$, $C_{i,y}$, and $C_{i,z}$ given in (19), (17), and (20), respectively;
@@
According to Definition 3, we have $\mathbf{x}_{j}^{t}+\boldsymbol{\vartheta}_{j}^{t}=\mathbf{x}_{j}^{\prime t}+\boldsymbol{\vartheta}_{j}^{\prime t}$, $y_{j}^{t}+\chi_{j}^{t}=y_{j}^{\prime t}+\chi_{j}^{\prime t}$, and $z_{j}^{t}+\zeta_{j}^{t}=z_{j}^{\prime t}+\zeta_{j}^{\prime t}$ for all $t\in\mathbb{N}$ and $j\in{\mathcal{N}_{i}}$. Since we assume that only the $k$th data point is different between $\mathcal{D}_{i,f}^{t}$ (resp. $\mathcal{D}_{i,g}^{t}$) and $\mathcal{D}_{i,f}^{\prime t}$ (resp. $\mathcal{D}_{i,g}^{\prime t}$), we have $\mathbf{x}_{i}^{p}=\mathbf{x}_{i}^{\prime p}$, $y_{i}^{p}=y_{i}^{\prime p}$, and $z_{i}^{p}=z_{i}^{\prime p}$ for all $p\leq k\leq t$. However, since the difference in loss functions kicks in at iteration $k$, i.e., $h(x,y;\varphi_{i}^{k})\neq h(x,y;\varphi_{i}^{\prime k})$ and $l(x;\xi_{i}^{k})\neq l(x;\xi_{i}^{\prime k})$, we have $\mathbf{x}_{i}^{t}\neq\mathbf{x}_{i}^{\prime t}$, $y_{i}^{t}\neq y_{i}^{\prime t}$, and $z_{i}^{t}\neq z_{i}^{\prime t}$ for all $t>k$. Hence, for the dynamics of $y_{i}^{t}$ in Algorithm 1, we have
@@
$\textstyle\quad+\lambda_{y}^{t}(g_{i}^{t}(R_{i}\mathbf{x}_{i}^{t})-g_{i}^{\prime t}(R_{i}\mathbf{x}_{i}^{\prime t}))\|_{1},$
@@
By using the relationship $\xi_{i}^{p}=\xi_{i}^{\prime p}$ for any $p\neq k$, the sensitivity $\Delta_{i,y}^{t+1}$ satisfies
$\textstyle\Delta_{i,y}^{t+1}$ $\textstyle\leq(1-|w_{ii}|)\Delta_{i,y}^{t}+\frac{\lambda_{y}^{t}}{t+1}\|l(R_{i}\mathbf{x}_{i}^{t};\xi_{i}^{k})-l(R_{i}\mathbf{x}_{i}^{\prime t};\xi_{i}^{\prime k})\|_{1}$
$\textstyle\quad+\frac{\sqrt{r}\lambda_{y}^{t}}{t+1}\sum_{p\neq k}^{t}\|l(R_{i}\mathbf{x}_{i}^{t};\xi_{i}^{p})-l(R_{i}\mathbf{x}_{i}^{\prime t};\xi_{i}^{p})\|_{2},$
@@
$\Delta_{i,y}^{t+1}\leq(1-|w_{ii}|)\Delta_{i,y}^{t}+L_{l}\sqrt{r}\lambda_{y}^{t}\Delta_{i,x}^{t}+\frac{2d_{i,l}\lambda_{y}^{t}}{t+1},$ (8)
where we have used the fact that the optimization variable $\mathbf{x}_{i}^{t}$ is always constrained in the compact set $\Omega$, i.e., $\max_{\mathbf{x}_{i}\in\Omega}\{\|l(R_{i}\mathbf{x}_{i};\xi_{i})\|_{1}\}\leq d_{i,l}$ holds for some $d_{i,l}>0$.
Furthermore, we use the projection inequality from Lemma 1 in [43] and the dynamics of $\mathbf{x}_{i}^{t}$ to obtain
$\textstyle\|\mathbf{x}_{i}^{t+1}-\mathbf{x}_{i}^{\prime t+1}\|_{1}\leq(1-|w_{ii}|)\|\mathbf{x}_{i}^{t}-\mathbf{x}_{i}^{\prime t}\|_{1}$
$\textstyle+\frac{\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla_{x}h(R_{i}\mathbf{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{p})-\nabla_{x}h(R_{i}\mathbf{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{\prime p})\|_{1}$
$\textstyle+\frac{\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla l(R_{i}\mathbf{x}_{i}^{t};\xi_{i}^{p})\tilde{z}_{i}^{t}-\nabla l(R_{i}\mathbf{x}_{i}^{\prime t};\xi_{i}^{\prime p})\tilde{z}_{i}^{\prime t}\|_{1}.$ (9)
@@
$\textstyle\frac{\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla_{x}h(R_{i}\mathbf{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{p})-\nabla_{x}h(R_{i}\mathbf{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{\prime p})\|_{1}$
@@
where in the derivation we have used the relationship $\|\tilde{y}_{i}^{t}-\tilde{y}_{i}^{\prime t}\|_{2}\leq\frac{1}{\lambda_{y}^{t}}(\|y_{i}^{t+1}-y_{i}^{\prime t+1}\|_{1}+\|y_{i}^{t}-y_{i}^{\prime t}\|_{1})$.